TANGENTS AND NORMALS
 at what points on the curve y = x^2 – 4x + 5 is the tangent perpendicular to the line 2y+x = 7

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Slope of line = -1/2
Slope of tangent = m
m(\frac{-1}{2}) = -1
m = 2
y = x^2 - 4x + 5
y' = 2x - 4
2x - 4 = 2
x = 3
Put in curve
y = 9 - 12 + 5
y = 2